OpenStudy (anonymous):

3 over 2 - 1 over x minus 4 = negative 2 over 2x minus 8 ?

OpenStudy (anonymous):

@alexwee123 another one! haha

OpenStudy (xguardians):

Lol can you use equation instead of writing it in letters , makes it easier to follow :P?

OpenStudy (anonymous):

\[\frac{ 3 }{ 2 }-\frac{ 1 }{ x-4 }=\frac{ -2 }{ 2x-8 }?\]

OpenStudy (alexwee123):

ok do you know what common denominators are ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

do you see how you could simplify 2x-8?

OpenStudy (anonymous):

*hint look at the denominators of your other fractions*

OpenStudy (alexwee123):

ok so what do you think the common demoniator between 3/2 and 1/(1-x) is?

OpenStudy (anonymous):

\[2\left( x-4 \right) ?\]

OpenStudy (alexwee123):

sorry 1/(x-4)*

OpenStudy (anonymous):

yes. now how do you change the denominators of your other fractions to get that same answer?

OpenStudy (anonymous):

This is where i'm stuck :/

OpenStudy (anonymous):

*hint you can multiply a fraction by another fraction that is equivalent to one*

OpenStudy (alexwee123):

yea awesome so if we simplify these two out we'd get